### v1

# Finding separatrices of dynamical flows with Deep Koopman Eigenfunctions

Abstract
Many natural systems, including neural circuits involved in decision making, can be modeled as high-dimensional dynamical systems with multiple stable states. While existing analytical tools primarily describe behavior near stable equilibria, characterizing separatrices – the manifolds that delineate boundaries between different basins of attraction – remains challenging, particularly in high-dimensional settings. Here, we introduce a numerical framework leveraging Koopman Theory combined with Deep Neural Networks to effectively characterize separatrices. Specifically, we approximate Koopman Eigenfunctions (KEFs) associated with real positive eigenvalues, which vanish precisely at the separatrices. Utilizing these scalar KEFs, optimization methods efficiently locate separatrices even in complex systems. We demonstrate our approach on synthetic benchmarks, ecological network models, and recurrent neural networks trained on neuroscience-inspired tasks. Moreover, we illustrate the practical utility of our method by designing optimal perturbations that can shift systems across separatrices, enabling predictions relevant to optogenetic stimulation experiments in neuroscience.
1 Introduction
Recurrent neural networks (RNNs) are widely used both in neuroscience, as models of circuit dynamics, and in machine learning, as powerful tools for sequential data processing [1, 2]. A key goal in neuroscience is to reverse-engineer these models in order to understand the underlying dynamical mechanisms [1, 2]. In particular, many cognitive tasks such as decision-making [3] and associative memory [4] can be modeled as multistable dynamical systems, where distinct decisions or memories correspond to different stable attractor states in phase space. Transitions between these attractors are governed by the geometry of the basins of attraction and, crucially, by the separatrix: the manifold that delineates the boundary between basins (Figure 1A).
A reverse-engineering method that has yielded significant insights about RNN computations involves finding approximate fixed points and linearising around them [5]. This involves minimizing a scalar function – the kinetic energy $q(x)=\|f(x)\|^{2}$ – to locate these points (Figure 1B). Once found, the linearisation of the dynamics at the fixed point can shed light on the mechanism of computations [6, 7, 8, 9, 10, 11, 12].
However, fixed points alone do not capture the global organization of multistable dynamics. Since inputs typically perturb the state in arbitrary directions, it is critical to know whether such perturbations cross the separatrix. To predict the effects of perturbations or design targeted interventions, one must characterize the separatrix itself.
Ideally, we would have a scalar function analogous to the kinetic energy – smooth, yet vanishing precisely on the separatrix (Figure 1C). This would allow gradient-based optimization to locate the decision boundary, help visualize the geometry of this manifold, and enable the design of optimal decision-changing perturbations (Figure 1A).
In this work, we propose a novel method to characterize separatrices in high-dimensional black-box dynamical systems by leveraging Koopman operator theory [13]. Specifically, we approximate scalar-valued Koopman eigenfunctions (KEFs) with positive real eigenvalues using deep neural networks. These eigenfunctions vanish precisely on the separatrix. Because they are scalar functions, gradient-based optimization can be used to efficiently trace out the separatrix, even in high dimensions.
We apply this framework to synthetic systems, ecological models, and RNNs trained on neuroscience-inspired tasks. In addition, we demonstrate that the learned KEFs can be used to design minimal perturbations that push the system across separatrices—a setting relevant to experimental protocols such as optogenetic stimulation.
Figure 1: (A) Phase-portrait of a 2D bistable system. The two attractors can signify different choices, and therefore the direction between them is called the choice direction. External input pushes the system across the separatrix letting it relax to the other attractor. The optimal perturbation has a different direction. (B) The kinetic energy vanishes at the fixed points (green ‘o’ stable and ‘+’ unstable) but does not reveal the full separatrix. (C) We aim to learn a scalar function $\psi(x)$ that vanishes precisely on the separatrix.
We summarise our main contributions:
  • We develop a tool to locate separatrices, the surfaces between basins of attraction in black-box multi-stable dynamical systems: a gap in the RNN reverse-engineering toolkit.
  • We demonstrate that KEFs with positive eigenvalues vanish precisely on the separatrix and can be trained using deep neural networks and a loss based on the Koopman PDE error.
  • We identify problems that may arise with the method. Mainly, two modes of degeneracy in the space of solutions to the PDE and propose effective regularization and training strategies to resolve them.
  • We show how the learned KEFs can be used to design minimal norm perturbations that shift the system across separatrices.
  • We empirically demonstrate the method on systems of increasing complexity: from 1D synthetic models to 64-dimensional trained RNNs, and an 11D ecological model.
2 Related Work
Our work builds on a growing body of literature at the intersection of Koopman operator theory, deep learning, and the analysis of dynamical systems, particularly in neuroscience and machine learning.
Koopman theory has recently been used to evaluate similarity between dynamical systems, both in neuroscience [14]—where it is applied to study the temporal structure of computation—and in machine learning, where it has been used to compare training dynamics across models [15]. These approaches typically analyze system-level behavior using dynamic mode decomposition [16, 17, 18] a finite-dimensional approximation of the Koopman Operator.
In parallel, deep learning methods have emerged as powerful tools for solving partial differential equations (PDEs). Notably, the Deep Ritz Method [19] and Deep Galerkin Method (DGM) [20] demonstrate how deep neural networks (DNNs) can approximate solutions to variational and differential problems. A related line of work uses physics-informed neural networks (PINNs), which incorporate known PDE constraints as part of the loss function during training [21].
Koopman-based embeddings have also been proposed as a tool for analyzing the internal dynamics of RNNs. In [22], the authors show that eigenvectors of finite-dimensional approximations of the Koopman operator can uncover task-relevant latent structure in RNNs. More generally, several works explore DNN-based approximations of Koopman operators for learning meaningful embeddings of nonlinear dynamics [23, 24, 25].
Finally, our approach is conceptually connected to work on the geometry of Koopman eigenfunctions themselves. In particular, [26] studies the level sets of KEFs and their relationship to isostables and isochrons in systems with stable fixed points. These constructions motivate our interest in identifying scalar functions that vanish on separatrices, as a means of understanding boundaries between basins of attraction.
3 Results
3.1 KEFs as Scalar Separatrix Indicators
We consider autonomous dynamical systems of the form:
$\dot{\boldsymbol{x}}=f(\boldsymbol{x}),\quad\boldsymbol{x}\in\mathcal{X}$ (1)
where the $\dot{\square}$ is shorthand for the time derivative $\frac{d}{dt}\square$ and $f:\mathcal{X}\to\mathcal{X}$ defines the dynamics on an $N$ dimensional state space $\mathcal{X}$.
Our goal is to construct a smooth scalar function $\psi$ that vanishes precisely on the separatrix between basins of attraction. A natural candidate for such a function is the inverse of the time it takes for a trajectory to reach an attractor. Near the separatrix, this time diverges, and hence the inverse goes to zero. Mathematically, we want an observable $\psi(\boldsymbol{x}(t))$ of the state of the system $\boldsymbol{x}$ that keeps growing as time evolves and $\boldsymbol{x}$ converges to an attractor of the system. This motivates a function $\psi(\boldsymbol{x})$ that, when $\boldsymbol{x}(t)$ is a trajectory of the dynamics, satisfies:
$\psi(\boldsymbol{x}(t))=\psi(\boldsymbol{x}(0))e^{\lambda t}.$ (2)
This is precisely the behavior of a Koopman eigenfunction (KEF) with eigenvalue $\lambda>0$. Note that Koopman eigenfunctions are usually introduced in a different manner, and the supplementary material shows the connection to our description.
A Koopman eigenfunction $\psi:\mathcal{X}\to\mathbb{R}$ satisfies:
$\frac{d}{dt}\bigg{(}\psi\big{(}\boldsymbol{x}(t)\big{)}\bigg{)}=\lambda\psi(\boldsymbol{x}(t))\quad\Leftrightarrow\quad\nabla\psi(\boldsymbol{x})\cdot f(\boldsymbol{x})=\lambda\psi(\boldsymbol{x}).$ (3)
This is the Koopman Partial differential equation (PDE) (see Appendix C for its relationship to the Koopman Operator). $\lambda$ relates to the timescale of $\psi$ and is an important hyperparameter of our method (Appendix I).
We approximate $\psi$ using a deep neural network (Appendix F) and train it by minimizing the Koopman PDE residual. Specifically, we define the loss:
$\mathcal{L}_{\text{PDE}}=\mathbb{E}_{\boldsymbol{x}\sim p(\boldsymbol{x})}\left[\nabla\psi(\boldsymbol{x})\cdot f(\boldsymbol{x})-\lambda\psi(\boldsymbol{x})\right]^{2},$ (4)
where $p(\boldsymbol{x})$ is a sampling distribution over the phase space [19, 20]. As with any eigenvalue problem, this loss admits the trivial solution $\psi\equiv 0$. To discourage such solutions, we introduce a shuffle-normalization loss where the two terms are sampled independently from the same distribution:
$\mathcal{L}_{\text{shuffle}}=\mathbb{E}_{\boldsymbol{x}\sim p(\boldsymbol{x}),\tilde{\boldsymbol{x}}\sim p(\boldsymbol{x})}\left[\nabla\psi(\boldsymbol{x})\cdot f(\boldsymbol{x})-\lambda\psi(\tilde{\boldsymbol{x}})\right]^{2},$ (5)
and optimize the ratio:
$\mathcal{L}_{\text{ratio}}=\frac{\mathcal{L}_{\text{PDE}}}{\mathcal{L}_{\text{shuffle}}}.$ (6)
Figure 2: Our method to approximate KEFs in three bistable systems. (A) A 1D system $\dot{x}=x-x^{3}$. The curve shows $f(x)$, and its fixed points in light green – ‘o’s stable and ‘x’s unstable. (B) The kinetic energy $q(x)$ of this system. (C) the true KEF (8) and it’s DNN approximation obtained by our method. (D,E,F) Damped Duffing Oscillator in 2D (G,H,I) 2-unit GRU [27] RNN trained on 1-bit flop flop (1BFF) [5] and our KEFs.
We train using stochastic gradient descent, where expectations are approximated by a batch of samples drawn from $p(\boldsymbol{x})$ and the shuffle corresponds to a random permutation of the samples in the batch (see Appendix G for details).
To illustrate the method, we start with an analytically solvable system in 1D (Figure 2A):
$\dot{x}=x-x^{3}$ (7)
The system has three fixed points, corresponding to minima of $q(x)$ (Figure 2B). A $\lambda=1$ KEF can be derived analytically (Appendix A)):
$\psi(x)=\frac{x}{\sqrt{|1-x^{2}|}}$ (8)
And the zero of this function corresponds to the unstable point, which serves as a separatrix in this 1D case. Figure 2C shows that the DNN approximates this function well, with the location of the zero (separatrix) being captured precisely.
We also apply the method to two 2D bistable systems: a 2D damped Duffing oscillator (Figure 2DEF), and a 2-unit GRU RNN trained on a one-bit flip-flop task (Figure 2GHI). In both cases, the system has two stable fixed points (green circles) and one unstable saddle (green crosses). Kinetic energy functions, shown for comparison, are minimized at the fixed points. In contrast, the learned $\lambda=1$ KEFs are zero on the separatrix (green contours).
3.2 Challenges and Solutions
While the examples above show cases where simple optimization leads to the separatrix, there are several crucial implementation details of our proposed methods. In particular, even a $\psi(x)$ that satisfies the Koopman PDE may fail to identify the true separatrix. This arises from known degeneracies in Koopman eigenfunctions, particularly in multistable or high-dimensional systems. To enable utilization of our tool, We describe two key failure modes and our strategies to resolve them, as summarized in Figure 3.
Degeneracy across basins.
A central issue stems from the compositional properties of Koopman eigenfunctions. Let $\psi_{1}(x)$ and $\psi_{2}(x)$ be eigenfunctions with eigenvalues $\lambda_{1}$ and $\lambda_{2}$. Then, their product is also a KEF:
$\nabla[\psi_{1}(x)\psi_{2}(x)]\cdot f(x)=(\lambda_{1}+\lambda_{2})\psi_{1}(x)\psi_{2}(x).$ (9)
In particular, consider a smooth KEF $\psi^{1}$ with $\lambda=1$ that vanishes only on the separatrix (e.g., as in Figure 2). Now, consider a piecewise-constant function $\psi^{0}$ with $\lambda=0$ that takes constant values within each basin and may be discontinuous at the separatrix. The product $\psi^{1}\psi^{0}$ remains a valid KEF with $\lambda=1$, but it can now be zero across entire basins—thereby destroying the separatrix structure we aim to capture (Figure 3 top).
We observe this behavior empirically in Appendix Figure 8, where independently initialized networks converge to different spurious solutions. To mitigate this, we introduce a balance regularization term that biases $\psi$ to have nonzero values in opposing basins, encouraging sign changes across the separatrix. Specifically, we define:
$\mathcal{L}_{\text{bal}}=\frac{(\mathbb{E}[\psi(x)])^{2}}{\text{Var}[\psi(x)]},$ (10)
and train using the combined loss $\mathcal{L}_{\text{ratio}}+\gamma_{\text{bal}}\mathcal{L}_{\text{bal}}$, where $\gamma_{\text{bal}}$ is a scalar hyperparameter.
Degeneracy in high dimensions.
In higher-dimensional systems, the Koopman PDE admits a family of valid KEFs that differ in their directional dependence. Consider a separable 2D system:
$\dot{x}=f_{1}(x),\quad\dot{y}=f_{2}(y).$ (11)
Solving the PDE for this system (appendix B) yields a family of KEFs parameterised by $\mu\in\mathbb{R}$:
$\psi(x,y)=A(x)^{\mu}B(y)^{1-\mu},$ (12)
where $A(x)$ and $B(y)$ are KEFs to the respective 1D problems. Crucially, when $\mu=1$, the eigenfunction depends only on $x$ and ignores $y$—capturing only the separatrix along $x=0$.
Figure 3 (bottom) illustrates this effect: different values of $\mu$ yield KEFs aligned with different separatrices. To address this, we train multiple KEFs $\{\psi_{i}(x)\}_{i=1}^{k}$ and combine them via the geometric mean:
$\psi_{\text{combined}}(x)=\left(\prod_{i=1}^{k}\psi_{i}(x)\right)^{1/k},$ (13)
which preserves the zero-level sets of all $\psi_{i}$.
To encourage each $\psi_{i}$ to capture different separatrix components, we propose sampling from multiple distributions centered somewhere on the separatrix. First, we obtain points on the separatrix by interpolating between fixed points and identifying the transition via binary search and numerical integration of the ODE (1). Around each such point $\beta_{j}$, we define a local distribution $\mathcal{N}(\beta_{j},\sigma_{j}^{2}I)$, using a range of scales $\{\sigma_{j}\}$ to span both fine and global structure. For each distribution, we compute losses $\mathcal{L}^{j}_{\text{ratio}}$ and $\mathcal{L}^{j}_{\text{bal}}$, and minimize the weighted sum:
$\mathcal{L}_{\text{total}}=\sum_{j=1}^{J}\mathcal{L}^{j}_{\text{ratio}}+\gamma_{\text{bal}}\mathcal{L}^{j}_{\text{bal}}.$ (14)
This procedure biases each KEF to resolve local separatrix geometry while retaining global consistency.
Figure 3: Top: In the presence of multiple basins, a KEF can collapse to zero within a single basin. This degeneracy is realised by multiplying the KEF with a piecewise constant KEF with $\lambda=0$ and invoking (9). This example corresponds to $\dot{x}=x-x^{3}$. We introduce a regularisation term (10) to encourage the mean value $\langle\psi\rangle\approx 0$. This encourages solutions with sign changes across basins. Bottom: In higher dimensions, degeneracy arises from directional ambiguity in solutions. We visualise the analytical solution (12) for $\dot{x}=x-x^{3};\dot{y}=y-y^{3}$. We address this by sampling from multiple local distributions around separatrix points and training an ensemble of KEFs.
3.3 Demonstrations
We demonstrate the applicability of the method on several qualitative examples.
3D GRU RNN Performing Two-Bit Flip Flop
We first demonstrate our method on a low-dimensional recurrent neural network trained to perform a two-bit flip flop (2BFF) task. Specifically, we use a 3-unit gated recurrent unit (GRU) network [27]. The trained network exhibits four stable fixed points (Figure 4), corresponding to different memory states of the task.
To overcome the degeneracies described in Figure 3, we adopt a targeted sampling strategy. We first identify points on the separatrix by interpolating between pairs of fixed points and performing binary search: at each step, we simulate the dynamics to determine basin membership and refine the search. Around these discovered separatrix points, we construct concentric isotropic Gaussian distributions, and sample from them to train on the loss $\mathcal{L}_{\text{total}}$, (14).
Two resulting KEF are shown in Figure 4 A,B). As expected, the KEFs vanish precisely along the separatrices. This result validates the ability of our method to recover boundary manifolds in neural dynamical systems, even in the presence of degeneracy. Once we know the separatrices, we can determine optimal perturbation directions (Figure 4C). Starting from a given initial condition (red star), we see that the same amplitude perturbation is sufficient to reach a different attractor when using the separatrix information, and insufficient when directed at the desired attractor. A more quantitative depiction of this effect is shown below in a higher-dimensional system.
Figure 4: Two-bit flip flop task in a 3-unit GRU. The system has 4 stable fixed points (light-green points). (A,B) Two KEFs obtained by our method. They complement each other as they each captures a separatrix along one direction. (C) Use of KEF to design minimal perturbations that push trajectories across the separatrix.
11D Ecological Dynamics
We next apply our method to a high-dimensional ecological model: a generalized Lotka–Volterra (gLV) system fit to genus-level abundance data from a mouse model of antibiotic-induced Clostridioides difficile infection (CDI) [28]. The system has five stable fixed points. Following [29] we focus our analysis to two of these fixed points representing healthy and diseased microbial states.
We optimize the KEF in the full 11-dimensional state space. For interpretability, we follow the projection approach of [29], visualizing the dynamics in the 2D plane spanned by the two chosen stable fixed points and the origin (see Figure 5). Although the KEF is trained entirely in the original 11-dimensional space, its zero level set (light green curve) aligns well with the true separatrix (orange line) computed using a grid of initial conditions in the 2D slice [29].
Figure 5: KEF approximation in a fitted 11D gLV model of CDI [28, 29]. Zero level set of the KEF aligns with the separatrix in a 2D projection plane.
This result demonstrates that our technique can be applied directly to real-world fitted models, without dimensionality reduction at training time.
Limit cycle separatrix
We test our method in a setting where there are no fixed points along the separatrix. We construct a system which oscillates at a fixed frequency ($\dot{\theta}=1$), but converges to one of two preferred amplitudes ($\dot{r}=(r-2)-(r-2)^{3}$). The system has three limit-cycles, two of them stable ($r=1,3$) and one unstable ($r=2$). In Figure 6B we visualise the flow, it’s kinetic energy and the limit cycles. The system has no fixed points, and thus fixed point analysis is futile. We utilize Radial basis function neural network [30] to parameterise the KEF (Appendix F).
We show that our approximation of the KEF recovers the separatrix at $r=2$ (Figure 6C).
Figure 6: Applying our method to a system of stable and unstable limit cycles, a system without any fixed points on the separatrix. (A) system equations. (B) kinetic energy, with dashed line for separatrix. (C) KEF from our method with zero level highlighted.
64D RNN Solving 1BFF
To test our method in a high-dimensional setting, we train a 64-unit vanilla RNN to solve a one-bit flip flop (1BFF) task. The resulting dynamics exhibit bistability between two memory states. As before, we identify a point on the separatrix via binary search along the line connecting the fixed point attractors, simulating the dynamics to classify basin membership at each step.
We then train the KEF network using samples drawn from isotropic Gaussian distributions centered at this separatrix point. Due to the high dimensionality, direct visualization of the KEF and the dynamics is not feasible. Instead, we devise a curve-based validation approach.
We construct multiple Hermite polynomial curves that interpolate between the two stable fixed points. The curvature of each curve is parameterized by a random vector, and each is defined by a parameter $\alpha\in[0,1]$, where $\alpha=0$ corresponds to one attractor and $\alpha=1$ to the other (see appendix E). Because each curve continuously connects the fixed points, it must cross the separatrix. Figure 7A shows a 2D PCA projection of several such Hermite curves. Crucially, the actual curves span the entire 64D space. We simulate dynamics from 100 points along each curve and determine their final basin to infer where each curve crosses the separatrix, forming a ground-truth reference.
Next, we evaluate the learned KEF along these same curves. Figure 7B shows KEF values along sample Hermite curves as a function of $\alpha$, with the zero crossing indicating our predicted separatrix. Figure 7C compares the $\alpha$-locations of the ground truth and the KEF-predicted separatrix points. We observe strong agreement, indicating that the learned KEF reliably tracks the separatrix in this high-dimensional system.
Finally, we demonstrate how the KEF can be used to design minimal perturbations that shift the state across the separatrix (similar to Figure 1A, Figure 4C). Given a base point $x_{\text{base}}$, we aim to solve:
$x^{*}=\operatorname*{arg\,min}_{x}\|x-x_{\text{base}}\|_{2}^{2}\quad\text{subject to}\quad|\psi(x)|=0,$ (15)
using the Adam optimizer [31] to minimize the relaxed loss:
$L(x)=|\psi(x)|+\gamma\|x-x_{\text{base}}\|_{2}^{2},$ (16)
with random initialization around $x_{\text{base}}$. Figure 7D compares the distance from $x_{\text{base}}$ to our optimal point versus separatrix points identified along Hermite curves and the direct line connecting the base point to the destination fixed point. Our method yields the shortest valid perturbation.
Figure 7: Validation of KEF approximation in a 64D RNN trained on 1BFF. (A) PCA projection of Hermite curves between fixed points coloured by true basin labels. (B) KEF values along three Hermite curves versus curve parameter $\alpha$, as well the true separatrix point along the curve. (C) Comparison between true and predicted separatrix positions along curves. (D) Distance from $x_{\text{base}}$ to perturbation targets. The KEF-guided solution yields the smallest perturbation crossing the separatrix.
4 Discussion
We presented a novel framework for identifying separatrices in high-dimensional, black-box dynamical systems using Koopman eigenfunctions (KEFs). This method is particularly useful for analyzing recurrent neural networks (RNNs), which are commonly used to model neural computations involving multiple stable states.
Prior efforts in reverse-engineering RNNs relied heavily on locating fixed points and linearizing dynamics locally [5, 6, 10, 12, 7, 8, 9, 11]. While powerful, these methods cannot directly capture global structures or predict system responses to large perturbations that cross basin boundaries. By directly approximating scalar-valued KEFs that vanish precisely on separatrices, our method complements and extends existing local linearization approaches. Practitioners can use our KEFs alongside fixed-point analysis to achieve a comprehensive understanding of the dynamical system’s landscape.
Our work also advances the application of Koopman operator theory to dynamical systems. Previous studies primarily utilized Koopman eigenfunctions to predict or control dynamics within a single basin of attraction [23, 32, 33, 34, 35]. Likewise, methods comparing dynamical systems to one another use the dynamic mode decomposition which does not always discern between different basins [14, 15, 36]. Such studies usually involve KEFs associated with negative eigenvalues ($\lambda$<0), which exhibit opposite behavior to ours: they explode at separatrices and approach zero at attractors. In contrast, we specifically targeted eigenfunctions associated with positive eigenvalues, ensuring their zeros correspond exactly to separatrices.
To help practitioners use our method, we highlight inherent challenges, such as degeneracy in the Koopman PDE, where multiple solutions exist. To overcome these, we introduced regularization strategies, notably a balance term ensuring eigenfunctions change sign across different basins. Additionally, ensemble training and targeted sampling methods were used to resolve directional ambiguities and ensure comprehensive coverage of separatrix geometry. These mitigations join existing work on KEF approximation [25, 24, 37] and enabled reliable identification of separatrices in diverse and high-dimensional systems.
Beyond the conceptual aspect of using positive KEFs to extract separatrices, and the qualitative manners in which this procedure can go wrong, there is also a computational aspect. While we did not elaborate on this here, it is important to note that finding the KEF is an iterative procedure in phase space. In contrast, it is possible to use grid searches and bisections to simulate the ODE in many locations to search for the separatrix [29]. The computational load of the latter approach scales with time, and requires to compute the dynamics associated with the same area in phase space many times. In contrast, solving the PDE is a form of dynamic programming that can be made more efficient. Specifically for the task of locating separatrices, critical slowing down can make solving the ODE computationally heavy.
While we demonstrated the applicability of our method to diverse scenarios, we do not provide theoretical guarantees linking the accuracy of the KEF approximation and that of the separatrix location. Furthermore, like techniques for finding fixed points [5, 38], our method requires knowing the dynamics in the entire phase space. In the future, extending this to trajectory-based methods [39, 40, 14] can facilitate the application of the method in neuroscience settings.
In conclusion, we hope that by focusing on separatrices, our method could inform intervention strategies in neuroscience, ecological or engineering systems, providing a general-purpose tool to predict and control transitions between stable states in complex dynamical landscapes.
Appendix A Analytical KEF derivation in 1D bistable system
We would like to find an analytical Koopman eigenfunction for the scalar dynamical system:
$\dot{x}$ $=x-x^{3}$ (17)
In the 1D case, the Koopman PDE (3) reduces to a first-order ordinary differential equation
$\frac{d\psi}{dx}f(x)$ $=\lambda\psi(x).$ (18)
With $f(x)=x-x^{3}$ and $\lambda=1$ we have:
$\psi^{\prime}(x)(x-x^{3})$ $=\psi(x)$ (19)
$\Rightarrow\quad\frac{\psi^{\prime}(x)}{\psi(x)}$ $=\frac{1}{x-x^{3}}$ (20)
To solve this integral we first simplify the integrand.
$x-x^{3}$ $=x(1-x^{2})=x(1-x)(1+x)$ (21)
So,
$\frac{1}{x(1-x)(1+x)}$ $=\frac{A}{x}+\frac{B}{1-x}+\frac{C}{1+x}$ (22)
Solving for $A$, $B$, $C$ yields $A=1$, $B=\frac{1}{2}$, $C=-\frac{1}{2}$.
Now we can integrate,
[…]
$=\log|x|-\frac{1}{2}\log|1-x|-\frac{1}{2}\log|1+x|+C$ (24)
$\log\psi(x)$ $=\log|x|-\frac{1}{2}\log|1-x|-\frac{1}{2}\log|1+x|+C$ (25)
$\Rightarrow\quad\psi(x)$ $=C^{\prime}\cdot\frac{|x|}{\sqrt{|1-x^{2}|}}$ (26)
To bring it into the form in the main text we use the product composition rule (9). We can multiply our solution by the $\operatorname{sign}(x)$ function which is a $\lambda=0$ eigenfunction because it remains constant in each basin (see Figure 3A). In other words, we flip the sign of our solution $\psi(x)\to-\psi(x)$ for $x<0$.
$\psi(x)$ $=C^{\prime}\frac{x}{\sqrt{|1-x^{2}|}}$ (27)
and this remains a KEF with $\lambda=1$.
Appendix B Eigenfunction Degeneracy in higher dimensions
Consider a separable 2D dynamical system:
$\dot{x}$ $=f_{x}(x),$ (28)
$\dot{y}$ $=f_{y}(y),$ (29)
which we write compactly as:
$\dot{\mathbf{x}}=\mathbf{f}(x,y)=\begin{bmatrix}f_{x}(x)\\ f_{y}(y)\end{bmatrix}.$ (30)
We seek a Koopman eigenfunction $\psi(x,y)$ satisfying:
$\nabla\psi\cdot\mathbf{f}(x,y)=\lambda\psi(x,y).$ (31)
Assume $\lambda=1$ and a separable form $\psi(x,y)=X(x)Y(y)$. Then:
$\frac{\partial\psi}{\partial x}$ $=X^{\prime}(x)Y(y),$ (32)
$\frac{\partial\psi}{\partial y}$ $=X(x)Y^{\prime}(y),$ (33)
$\nabla\psi\cdot\mathbf{f}$ $=X^{\prime}(x)Y(y)f_{x}(x)+X(x)Y^{\prime}(y)f_{y}(y)=X(x)Y(y).$ (34)
Dividing both sides by $X(x)Y(y)$ gives:
$\frac{X^{\prime}(x)}{X(x)}f_{x}(x)+\frac{Y^{\prime}(y)}{Y(y)}f_{y}(y)=1.$ (35)
The above equation requires that the sum of the above two terms, which each depend on different variables must be 1 for all $x$, $y$. It follows that each term is also a constant function.
$\frac{X^{\prime}(x)}{X(x)}f_{x}(x)$ $=\mu,$ (36)
$\frac{Y^{\prime}(y)}{Y(y)}f_{y}(y)$ $=1-\mu,$ (37)
for an arbitrary constant $\mu\in\mathbb{R}$.
Define the antiderivatives:
[…]
Then the logarithms of the separated components are:
$\log X(x)$ $=\mu A(x)\quad\Rightarrow\quad X(x)=\left(e^{A(x)}\right)^{\mu},$ (40)
$\log Y(y)$ $=(1-\mu)B(y)\quad\Rightarrow\quad Y(y)=\left(e^{B(y)}\right)^{1-\mu}.$ (41)
Thus, the general separable Koopman eigenfunction is:
$\psi(x,y)=\left(e^{A(x)}\right)^{\mu}\cdot\left(e^{B(y)}\right)^{1-\mu}.$ (42)
Appendix C Relation of our definition to the Koopman Operator
In the main text, we introduced Koopman eigenfunctions as scalar functions $\psi:\mathcal{X}\to\mathbb{R}$ that evolve exponentially along trajectories $\boldsymbol{x}(t)\in\mathcal{X}$ of a dynamical system $\dot{\boldsymbol{x}}=f(\boldsymbol{x})$:
$\frac{d}{dt}\psi(\boldsymbol{x}(t))=\lambda\psi(\boldsymbol{x}(t)).$ (43)
Here, we clarify the origin of this equation by defining the Koopman operator, linking our approach to the broader theory.
Let $g:\mathcal{X}\to\mathbb{R}$ be a real-valued function of the system state—commonly referred to as an observable. The collection of such observables forms an infinite-dimensional function space, typically a Hilbert space once equipped with an inner product $\langle g,g^{\prime}\rangle$. The Koopman operator acts linearly on this space.
For a continuous-time system, the Koopman operator $\mathcal{K}_{\tau}$ evolves observables according to the flow map $F_{\tau}:\mathcal{X}\to\mathcal{X}$, which advances the state forward by time $\tau$:
$(\mathcal{K}_{\tau}g)(\boldsymbol{x}(t))=g(F_{\tau}(\boldsymbol{x}(t)))=g(\boldsymbol{x}(t+\tau)).$ (44)
The infinitesimal generator of the Koopman semigroup $\{\mathcal{K}_{\tau}\}_{\tau\geq 0}$, often denoted simply as $\mathcal{K}$, is defined as:
$\mathcal{K}g:=\lim_{\tau\to 0}\frac{\mathcal{K}_{\tau}g-g}{\tau}=\lim_{\tau\to 0}\frac{g(F_{\tau}(\boldsymbol{x}))-g(\boldsymbol{x})}{\tau}.$ (45)
When evaluated along a trajectory $\boldsymbol{x}(t)$, this yields:
$\mathcal{K}g(\boldsymbol{x}(t))=\lim_{\tau\to 0}\frac{g(\boldsymbol{x}(t+\tau))-g(\boldsymbol{x}(t))}{\tau}=\frac{d}{dt}g(\boldsymbol{x}(t))=\nabla g\cdot\boldsymbol{\dot{x}}(t)=\nabla g\cdot f\big{(}\boldsymbol{x}(t)\big{)}.$ (46)
This operator is also known as the Lie derivative of $g$ along the vector field $f$.
Thus, an eigenfunction $\psi$ of $\mathcal{K}$ satisfying
$\mathcal{K}\psi=\lambda\psi$ (47)
recovers the Koopman eigenfunction equation (3) used in the main text.
Appendix D KEF degeneracy in randomly initialised DNN solutions
Main text Figure 3 illustrates challenges arising due to the degeneracy of the Koopman PDE (3). In Figure 8, we train several DNNs on a 2-unit GRU trained on the 2BFF. Each DNN is independently initialised and trained on a single distribution without the balance regularisation term $\mathcal{L}_{\text{bal}}$, i.e., $\gamma_{\text{bal}}=0$. The resulting KEF approximations exhibit the same modes of degeneracy - zero on certain basins as well as vertical and horizontal variants.
Figure 8: Many KEFs of to for 2 bit flip flop in 2D
Appendix E Curve-based validation approach
In high dimension, we cannot visualize the entire phase space to check whether zeros of the KEF coincide with the separatrix. Instead, we generate a family of smooth curves that connect two attractors, and hence must pass through a separatrix. In the 64D GRU flip flop example, the two attractors are the two stable fixed points $x,y\in\mathbb{R}^{N}$. We use cubic Hermite interpolation with randomized tangent vectors at the endpoints. Each curve is defined by:
$H(\alpha)$ $=h_{00}(\alpha)\,x+h_{10}(\alpha)\,m_{x}+h_{01}(\alpha)\,y+h_{11}(\alpha)\,m_{y},\quad\alpha\in[0,1]$ (48)
where the Hermite basis functions are:
$h_{00}(\alpha)$ $=2\alpha^{3}-3\alpha^{2}+1,$ (49)
$h_{01}(\alpha)$ $=-2\alpha^{3}+3\alpha^{2},$ (50)
$h_{10}(\alpha)$ $=\alpha^{3}-2\alpha^{2}+\alpha,$ (51)
$h_{11}(\alpha)$ $=\alpha^{3}-\alpha^{2}.$ (52)
Notice that $H(0)=x$ and $H(1)=y$.
The tangent vectors $m_{x}$ and $m_{y}$ are initialized as $y-x$ and perturbed with Gaussian noise:
$m_{x}=(y-x)+\epsilon_{x},\quad m_{y}=(y-x)+\epsilon_{y},\quad\epsilon_{x},\epsilon_{y}\sim\mathcal{N}(0,\sigma^{2}I).$ (53)
We sample multiple such curves with independently drawn perturbations. This produces a family of curves that interpolate between $x$ and $y$, while varying in geometry, enabling randomized exploration of intermediate regions in state space. Crucially, the curves are not limited to the manifold spanned by the attractors, but extend to all dimensions (controlled by $\sigma$. Optional constraints (e.g., non-negativity) can be imposed by rejecting any curve that violates them. For each such curve, we both evaluate the KEF and simulate the ODE to determine the position of the separatrix.
Appendix F Neural network architectures
In most of the demonstrations we use a ResNet architecture [41] with a tanh activation function.
Let the input to the network be $\mathbf{x}_{\text{in}}\in\mathbb{R}^{d_{\text{in}}}$ and define the hidden layer activations $\mathbf{x}_{\text{hid}}\in\mathbb{R}^{d_{\text{hid}}}$, output dimension $\mathbf{x}_{\text{out}}\in\mathbb{R}^{d_{\text{out}}}$, and number of layers $L$.
The network receives inputs at the first layer $\mathbf{x}^{(0)}=\text{Pad}(\mathbf{x}_{\text{in}}$), where Pad appends zeroes to the input as we always choose $d_{\text{hid}}>d_{\text{in}}$. The network then transforms the input at each layer $l$,
$\mathbf{x}^{(\ell+1)}=\mathbf{x}^{(\ell)}+\tanh\left(W^{(\ell)}\mathbf{x}^{(\ell)}+\mathbf{b}^{(\ell)}\right),\quad\ell=1,\dots,L-1,$ (54)
where $W^{(l)}\in\mathbb{R}^{d_{\text{hid}}\times d_{\text{hid}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{\text{hid}}}$
The output is the final Residual layer activation after applying another linear layer to match the desired $d_{\text{out}}$:
$\mathbf{x}_{\text{out}}=W^{\text{out}}\mathbf{x}^{(L)}+\mathbf{b}^{\text{out}},$ (55)
where $W^{\text{out}}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{hid}}}$ and $\mathbf{b}^{\text{out}}\in\mathbb{R}^{d_{\text{out}}}$. During optimization, gradients $\nabla\theta\mathcal{L}_{\text{total}}$ are computed for all parameters $\theta=(W^{(1:L-1)},\textbf{b}^{(1:L-1)},W^{\text{out}},\textbf{b}^{\text{out}})$.
Choices for each system
For the results in Figures 2, 4, 7, we use $L=20$, $d_{\text{hid}}=400$. $d_{\text{out}}=1$ and $d_{\text{in}}=N$ the dimension of the dynamical system. For the gLV system in Figure 5 we use $L=25$ and $d_{\text{hid}}=1000$.
Radial Basis Function (RBF) Layer
For the limit cycles example in Figure 6 we use a single Radial Basis Function layer [30].
Given an input $\mathbf{x}\in\mathbb{R}^{d_{\text{in}}}$, the RBF layer maps it to an output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ through a set of $M$ radial basis functions, each centered at $\mathbf{c}_{i}\in\mathbb{R}^{d_{\text{in}}}$, with a shape parameter $\varepsilon_{i}>0$ and linear combination weights $a_{ij}\in\mathbb{R}$.
To compute RBF activations for $i=1,\dots,M$, we define the scaled radial distance:
$s_{i}(\mathbf{x})=\varepsilon_{i}\cdot\left\|\mathbf{x}-\mathbf{c}_{i}\right\|,$ (56)
and then apply a gaussian radial basis function
$\varphi_{i}(\mathbf{x})=\exp(-s_{i}(\mathbf{x})^{2}).$ (57)
The final output is a linear combination of the basis activations:
$y_{j}(\mathbf{x})=\sum_{i=1}^{M}a_{ji}\cdot\varphi_{i}(\mathbf{x}),\quad j=1,\dots,d_{\text{out}}$ (58)
We use $M=300$, and $d_{\text{out}}=1$. During optimization, gradients $\nabla\theta\mathcal{L}_{\text{total}}$ are computed for all parameters $\theta=(\{a_{ji}\},\{\mathbf{c}_{i}\},\{\varepsilon_{i}\})$.
Appendix G Optimisation
We minimise the total loss:
$\mathcal{L}_{\text{total}}=\sum_{j=1}^{J}\mathcal{L}^{j}_{\text{ratio}}+\gamma_{\text{bal}}\mathcal{L}^{j}_{\text{bal}}.$ (59)
where $j$ corresponds to the $j^{\text{th}}$ sampling distribution (see main text section 3.2). $B$ $N$-dimensional points in the state space $\mathcal{X}$ are sampled from each distribution $\boldsymbol{x}^{j}\sim p_{j}(\boldsymbol{x})$. The ratio loss is the Koopman PDE error, normalised by a sample-shuffled version:
$\mathcal{L}_{\text{ratio}}^{j}$ $=\frac{\sum_{i=1}^{B}(\text{LHS}_{i}^{j}-\text{RHS}_{i}^{j})^{2}}{\sum_{i=1}^{B}(\text{LHS}_{i}^{j}-\text{RHS}_{\text{perm}(i)}^{j})^{2}}$ (60)
$\text{LHS}_{i}^{j}$ $=\nabla\psi(\boldsymbol{x}_{i}^{j})\cdot f(\boldsymbol{x}_{i}^{j})$ left-hand-side of the Koopman PDE (3) (61)
$\text{RHS}_{i}^{j}$ $=\lambda\psi(\boldsymbol{x}_{i}^{j})$ right-hand-side of the Koopman PDE (3) (62)
where $\text{perm}(i)$ is a random permutation of the numbers $1,2,\dots,B$ sampled during each training iteration.
The balance regularisation loss is the squared mean of the KEF values divided by their variance:
$\mathcal{L}_{\text{bal}}^{j}$ $=\frac{(\bar{\psi}^{j})^{2}}{\frac{1}{B}\sum_{i=1}^{B}(\psi(\boldsymbol{x}_{i}^{j})-\bar{\psi}_{j})^{2}},$ (63)
$\bar{\psi}^{j}$ $=\frac{1}{B}\sum_{i=1}^{B}\psi(\boldsymbol{x}_{i}^{j}).$ (64)
In general we set $\gamma_{\text{bal}}=0.05$. For the limit cycles Figure 6 we set $\gamma_{\text{bal}}=0$.
We compute $\nabla\psi(x)$ using Pytorch’s torch.autograd.grad, specifying create_graph=True, since we differentiate through this a second time to compute the gradients $\nabla_{\theta}\mathcal{L}_{\text{total}}$ with respect to the neural network parameters $\theta$.
We use the Adam optimiser [31] with learning rate $10^{-4}$ and l2 normalisation $10^{-5}$. We use $B=1000$ and train for $1000$ iterations.
Only in the case of the 11D gLV, Figure 5 we use $B=5000$ and train for $5000$ iterations.
A summary of all hyperparameters is provided in Table 1.
Table 1: Algorithm details and hyperparameters for various systems. System dimensionality $N$, Koopman eigenvalue $\lambda$, balance regularisation weight $\gamma_{\text{bal}}$, batch-size $B$, training iterations $T$, learning rate $\eta$, ResNet depth $L$ and width $d_{\text{hid}}$, number of Radial Basis Functions $M$.
Dynamical System $N$ $\lambda$ $\gamma_{\text{bal}}$ $B$ $T$ $\eta$ $L$ $d_{\text{hid}}$ $M$
Bistable 1D $1$ $1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
Damped Duffing oscillator $2$ $1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
1BFF, 2D GRU $2$ $1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
2BFF, 3D GRU $3$ $0.2$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
1BFF, 64D $64$ $0.1$ $5\times 10^{-2}$ $1000$ $1000$ $10^{-4}$ $20$ $400$
Two Limit Cycles $2$ $1$ 00 $1000$ $1000$ $10^{-4}$ $300$
Ecology gLV [28, 29] $11$ $0.1$ $5\times 10^{-2}$ $5000$ $5000$ $10^{-4}$ $25$ $1000$
Appendix H Computational Resources and run time
We ran all experiments on a system with four GeForce GTX 1080 GPUs with 10 Gbps of memory each.
All the 2D systems take 1-5 minutes to train the KEFs. The 11D gLV takes up to 20 minutes. The 64D RNN performing 1BFF takes 5 minutes to train the KEF.
Appendix I Choice of eigenvalue for numerics
In the main text we look for approximations to the Koopman PDE (3) for a real positive eigenvalue $\lambda$. What should the value of $\lambda$ be? It is known that products of KEFs are KEFs themselves with different eigenvalues. In particular, for a KEF $\psi$ with eigenvalue $\lambda$, we see that:
$\nabla\left[\psi(x)^{\alpha}\right]\cdot f(x)$ $=\alpha\psi(x)^{\alpha-1}\nabla\psi(x)\cdot f(x)$ (65)
$=\alpha\lambda\psi(x)^{\alpha}$ (66)
Therefore, $\psi(x)^{\alpha}$ is also a Koopman eigenfunction, with eigenvalue $\alpha\lambda$. This translates to changes in the shape of the KEF, i.e., the sharpness of the peaks, while maintaining the position of the zeroes.
In practice the choice of $\lambda$ affects training convergence, and it is therefore an important hyperparameter in the optimisation procedure (see Figure 9). We attribute this to the time scale of interest in the system $\boldsymbol{\dot{x}}=f(\boldsymbol{x})$, and differences in the propagation of gradients for different $\lambda$.
Figure 9: Training convergence as a function of eigenvalue $\lambda$, evaluated by normalised PDE error $\mathcal{L}_{\text{ratio}}$ 6 for two systems: LEFT, 1D bistable system $\dot{x}=x-x^{3}$ (see Figure 2) and 2BFF GRU 3D (see Figure 4).